# Model-agnostic Coreset Selection via LLM-based Concept Bottlenecks

### v1

@@ 5.4 Visualizing samples with their concept-based scores
Here we show the advantage of our concept-based score for assessing the sample’s difficulty and how using concepts aids its interpretability. We start by visualizing the easiest and the most challenging images (per class) for CIFAR-10/100. In Fig. 2, the top row shows the images with the highest concept-based scores (easiest) and the bottom shows the images with the lowest scores (challenging) for a subset of classes in CIFAR-10/100. As observed the easiest images are typical images associated with the label where as the challenging images are confusing (and even potentially mislabeled) as they look like images from a different class. For example, some challenging images in the class “boy” from CIFAR-100 are actually images of a baby which is also a class in CIFAR-100. Similarly, some challenging images from the class “cat” in CIFAR-10 look like images of a dog. More examples of such images are presented in Fig. 3 in App. C.1. Since the challenging examples are confusing for humans too, any ML model will find them hard to learn as well. This shows that our score is well aligned with the human-perceived difficulty of the samples.
@@
Next, we demonstrate why certain samples get low/high concept-based scores in our approach by extracting attributes specific to these images using LLaVA (note that these attributes are different from the per-class concepts used in the concept bottleneck). To generate these, we prompt LLavA to produce concepts using both the sample image and its class label (see image-level concept extraction in App. B). These image-level attributes are shown in the boxes in Fig. 2. As observed in Fig. 2, image-level attributes provided by LLaVA are related to the class label for easy images whereas they are unrelated for challenging images. For example, attributes provided by LLaVA for the challenging images of “airplane” align more with those of a ship (both of which are classes in CIFAR-10), and concepts provided for challenging images of “bridges” align more with those of castles (both of which are classes in CIFAR-100). Since our score also assigns a small value for these images, our concept-based score in Eq. 4 can correctly capture when the visual information in the sample is not aligned with the associated label of the sample and vice-versa. Thus, explaining why certain examples should be included/excluded from the coreset in a human-understandable way and at a dataset level independent of the downstream model.
@@ 5.5 Ablation studies
Here, we present an ablation to study the effect of different methods for concept generation and the effect of keeping different numbers of concepts per class ($k$) in the bottleneck.
@@
Comparison of different techniques to generate concepts via LLaVA. Table 4 shows how the performance of models trained on our coresets change when different methodologies are used to generate the concept sets. Since our method uses LLaVA (Liu et al., 2023b), which is a vision language model, we compare the performance of models trained on the random subsets and coresets obtained using class-wise concepts (only textual information) and concepts extracted using both visual and textual information. For concepts generated using only textual information, we consider two alternatives, namely class-wise attributes (CW-A) and class-wise descriptions (CW-D). While CW-A considers concepts formed by a single or a few words, CW-D consists of longer, more descriptive concepts (eg., a descriptive concept for the class butterfly is “a beautiful insect with colorful wings”). For CW-D, we use a subset of $k$ concepts provided by (Yang et al., 2023b), generated via the GPT-3 model. Our results show that CW-A performs better than CW-D for both the standard and label-free CS problems. Thus, we use attribute-level concepts in our work.
Table 4: Coreset performance on CIFAR-100 for $90\%$ pruning rate for concepts generated using different methods. A random subset of CIFAR-100 achieves an accuracy of $44.76_{\tiny\pm 1.58}$ at this rate. (Class-wise attributes approach is used for in rest of the paper.)
Concept generation methods
Prompt uses only class label Prompt uses class label + image(s)
class-wise attributes class-wise descriptions class-wise one-shot image attributes image-wise attributes
Ours $51.85_{\tiny\pm 0.51}$ $51.05_{\tiny\pm 0.71}$ $51.68_{\tiny\pm 0.45}$ $52.47_{\tiny\pm 0.39}$
Ours-LF $49.91_{\tiny\pm 0.96}$ $49.85_{\tiny\pm 0.83}$ $50.22_{\tiny\pm 0.16}$ $51.05_{\tiny\pm 0.93}$
@@
Next, for generating concepts using both visual and textual information, we consider two alternatives. The first is a class-wise one-shot image attribute approach where we first cluster all images of a class in the embedding space of the CLIP’s visual encoder and identify the image whose embedding is the closest to the cluster center (for the label-free setting we use the pseudo-labels of the images during clustering ), then we prompt LLaVA to generate attributes using this single image and the class name. Once generated we use $k$ discriminative concepts to form the bottleneck. The second is image-wise attribute approach, where we use each image in the training set and prompt LLaVA to generate per image attributes describing the image. Once generated we sort the concepts based on their frequency of occurrence in a class and use the most frequently occurring discriminative concepts to form the bottleneck. While the image-level concepts lead to the best coresets, it is slow and costly to prompt LLaVA to generate attributes for all the images in a large dataset such as Imagenet. For CIFAR-100, this process took about nine hours to complete (in comparison CW-A can be extracted in 5 minutes for CIFAR-100, without parallel computation) which is very costly compared to the small performance gains it provides over other approaches. Lastly, while the one-shot approach is better than CW-A in most cases, the additional step of clustering can be costly for larger datasets such as Imagenet. Based on these results, we used CW-A for concept generation using LLaVA.
Effect of $k$. In Table 5, we show how the number of concepts extracted per class label affects the selection of coresets. Once the list of attribute-level concepts is generated by LLaVA, we can select $k$ concepts per class either randomly or choose concepts unique to a class (discriminative). Our results show that using even $k$=$1$ is sufficient to surpass the performance using a random subset (Random baseline). This performance increases when we keep discriminative concepts in our concept bottleneck, with $k=5$ achieving the best results. While the size of the concept bottleneck need not be very large to find good coresets, it is helpful to take a sample’s visual similarity with a set of concepts rather than a single concept per class. Thus, we selected $5$ discriminative concepts per class to form the concept bottleneck.
@@
Table 5: Effect of the number of concepts $k$ (per class) and the method of selecting $k$ attribute-level concepts from the concepts generated by LLaVA vs. the accuracy of models trained on the coresets of CIFAR-100 at $90\%$ pruning rate. A random subset of CIFAR-100 achieves an accuracy of $44.76_{\tiny\pm 1.58}$ at this rate.
@@
Random concepts $48.78_{\tiny\pm 0.96}$ $50.15_{\tiny\pm 1.64}$ $50.39_{\tiny\pm 0.72}$
Discriminative $51.42_{\tiny\pm 0.18}$ $\mathbf{51.58}_{\tiny\pm 0.51}$ $51.22_{\tiny\pm 0.72}$
@@ 6 Conclusion
CS finds representative samples from the training data, training models on which leads to models with accuracy similar to the models trained on the entire dataset. In this work, we proposed a score based on concept bottlenecks that allow us to gauge the difficulty of a sample in terms of human-understandable concepts and is independent of the downstream model. Our experiments show that training downstream models on coresets selected using our score and a stratified sampling approach leads to better performance than random subsets and achieves accuracy similar to or better than the SOTA approaches based on training dynamics of the downstream model, for both the standard and label-free CS problem. Moreover, our score provides an intuitive explanation of a sample’s difficulty independent of the downstream model and leads to coresets that achieve high accuracy for various model architectures.
Limitations: In our work, the concept extraction from LLM is treated as a pre-processing step which is independent of the data difficulty score computation and coreset selection. However, LLMs can produce very noisy and non-discriminative concepts for various classes (eg., a concept “blue” can be associated with several classes in the Imagenet dataset), leading to poor concept-similarity scores. Moreover, while class-wise concept extraction is efficient, image-level concepts could be much more informative for a sample’s difficulty. Thus, improving the efficiency of concept extraction and tuning the prompt for LLMs/VLMs to incorporate the feedback from score computation or CS can help generate better concept bottlenecks leading to a better estimate of a sample’s difficulty. While these are important research directions we leave them for future work.
@@
Appendix A Additional related work on concept-based interpretability
Interpretability methods can be broadly classified as post-hoc methods (do not impose any model constraints) or by-design methods. Post-hoc methods include Gradient-weighted Class Activation Mapping approaches (Bau et al., 2017; Selvaraju et al., 2017; Mu & Andreas, 2020; Hernandez et al., 2021) that trace network gradients to identify the input areas that guide predictions and Explanation Generation methods (Singh et al., 2023; Nishida et al., 2022; Kim et al., 2018; Hendricks et al., 2016) that require models to produce explanations for visual tasks by conditioning their predictions on captioning models or incorporating visual evidence to ground explanations (Hendricks et al., 2018; Park et al., 2018). Interpretable-by-design methods, such as Prototype methods, optimize a metric space where classifications are based on distances to class prototypes, identifying important input regions but often obscuring their semantic content (Nauta et al., 2021; Chen et al., 2019; Snell et al., 2017; Satorras & Estrach, 2018; Vinyals et al., 2016).
Concept Bottleneck Models (CBMs) are a part of interpretable-by-design approaches that use human-understandable attributes as an intermediate layer for predictions. A recent advancement, Computational Derivation Learning (CompDL), utilizes a CBM architecture by applying a linear layer over CLIP scores between expert-designed concepts and images, improving evaluation of how well CLIP grounds concepts (Yun et al., 2022). Post-hoc Concept Bottleneck Models (PCBMs) were recently proposed to ease the requirement of CBMs to rely on costly concept annotations and improve their accuracy compared to end-to-end models. However, PCBMs are limited by the coverage of knowledge bases, making them unsuitable for large-scale or domain-specific tasks or fine-grained classification, and their residual predictors can undermine interpretability by blending CBMs with end-to-end models.
High-level semantic-driven descriptions are also used to guide data augmentation to build an informative set (Wickramanayake et al., 2021) to make model training efficient with a good enough training set. Prior works use external knowledge bases to obtain these textual semantic concepts to guide vision models (Bujwid & Sullivan, 2021; Kil & Chao, 2021; Roth et al., 2022; Shen et al., 2022). Thus, the use of concepts has been shown to improve interpretability in various domains. However, to the best of our knowledge, we are the first ones to propose a concept-based score for the CS problem and show its competitiveness to SOTA model training dynamics-dependent approaches.
@@ Appendix B Details for concept set generation
Alternative VLMs for Concept Generation: We leverage LLaVA as our choice of VLM for concept generation, however in Table 4, we also compare against concepts extracted from GPT (column 2) (Yan et al., 2023). We see comparable performance with those extracted from LLaVA. Moreover, LLaVA is an open-source model whereas GPT prompting is not. We also experimented with retrieving concepts via SpLiCE (Bhalla et al., 2024). However, a major limitation of SpLiCE is that similar to image-level concepts it is a costly approach. SpLiCE uses a linear optimization for sparse concept decomposition and can take up to 3 hours for $50,000$ images which is slower than generating class-level concepts from LLaVA.
@@ C.1 Visualizing easy/challenging samples based on concept-based score
C.2 Algorithm for stratified sampling using CCS (Zheng et al., 2022)

### v2

@@ 5.3 Visualizing easy and challenging samples based on our concept-based score
Here we show the advantage of our concept-based score for assessing the sample’s difficulty and how using concepts aids its interpretability. We start by visualizing the easiest and the most challenging images (per class) for CIFAR-10/100. In Fig. 3, the top row shows the images with the highest concept-based scores (easiest) and the bottom shows the images with the lowest scores (challenging) for a subset of classes in CIFAR-10/100. As observed the easiest images are typical images associated with the label where as the challenging images are confusing (and even potentially mislabeled) as they look like images from a different class. For example, some challenging images in the class “boy” from CIFAR-100 are actually images of a baby which is also a class in CIFAR-100. Similarly, some challenging images from the class “cat” in CIFAR-10 look like images of a dog. More examples of such images are presented in Fig. 4 in App. C.1. Since the challenging examples seem to be confusing, it shows that our score can identify examples that are ambiguous or mislabeled. Such samples may be hard for some ML models to learn and could force them to rely on spurious features, potentially hurting generalization. The ability of our score to identify such samples without the downstream model demonstrates its effectiveness for ranking the samples for CS.
@@
Next, we demonstrate why certain samples get low/high concept-based scores in our approach by extracting concepts specific to these images using LLaVA (note that these concepts are different from the per-class concepts used in the concept bottleneck). To generate these, we prompt LLaVA to produce concepts using both the sample image and its class label (see image-level concept extraction in App. B). These image-level concepts are shown in the boxes in Fig. 3. As observed in Fig. 3, image-level concepts provided by LLaVA are related to the class label for easy images whereas they are unrelated for challenging images. For example, attributes provided by LLaVA for the challenging images of “airplane” align more with those of ship (both which are classes in CIFAR-10), and concepts provided for challenging images of “bridges” align more with those of castles (both of which are classes in CIFAR-100). Since our concept-based score in Eq. 4 assigns a small value for these images, we see that it correctly captures when a sample’s visual information is not aligned with the sample’s associated label and vice-versa. Thus, explaining why certain examples should be included/excluded from the coreset in an interpretable way independent of the downstream model.
@@ 5.4 Ablation studies
Here, we present ablation studies to evaluate various component of our approach. Specifically, we evaluate the effect of different 1) methods and LLMs used for concept generation, 2) number of concepts per class ($k$) in the bottleneck, 3) CLIP backbones for visual/concept similarity, 4) number of iterations ($T$) for training the concept bottleneck layer. Here we use CIFAR-100 with a $90\%$ pruning ratio and train a ResNet-18 on the selected coresets. Lastly, we also evaluated the effectiveness of our method for CS on tasks beyond object recognition using medical and affective computing tasks.
@@
Comparison of different techniques to generate concepts via LLaVA. Tables 2 and 5 shows how the performance of models trained on our coresets change when different method are used to generate the concept sets (Fig. 2 block 1). Since our method uses LLaVA liu2023llava , which is a VLM, we compare the performance of models trained on the random subsets and coresets obtained using class-wise concepts (only textual information) and concepts extracted using both visual and textual information.
Table 4: Effect of using concepts obtained using only textual information (attributes vs. descriptions) to form the concept bottleneck layer  [^2] Cf. A random subset of CIFAR-100 achieves an accuracy of $44.76_{\pm 1.58}$ at 90% pruning rate..
class-wise attributes class-wise descriptions
Ours $51.85_{\scriptscriptstyle\pm 0.51}$ $51.05_{\scriptscriptstyle\pm 0.71}$
Ours-LF $49.91_{\scriptscriptstyle\pm 0.96}$ $49.85_{\scriptscriptstyle\pm 0.83}$
For concepts generated using only textual information, we consider two alternatives, namely class-wise attributes (CW-A) and class-wise descriptions (CW-D). While CW-A considers concepts formed by a single or a few words, CW-D consists of longer, more descriptive concepts (eg., a descriptive concept for the class butterfly is “a beautiful insect with colorful wings"). For CW-D, we use a subset of $k$ concepts provided by yang2023language , generated via the GPT-3 model. Our results show that CW-A performs better than CW-D for both the standard/label-free CS problems. Thus, we used CW-A for all our experiments.
Table 5: Effect of using concepts generated by using both visual and textual information to form the concept bottleneck layer footnotemark: .
class-wise one-shot image attributes image-wise attributes
Ours $51.68_{\scriptscriptstyle\pm 0.45}$ $52.47_{\scriptscriptstyle\pm 0.39}$
Ours-LF $50.22_{\scriptscriptstyle\pm 0.16}$ $51.05_{\scriptscriptstyle\pm 0.93}$
@@
Next, for generating concepts using both visual and textual information, we consider two alternatives. The first is a class-wise one-shot image attribute approach where we first cluster all images of a class in the embedding space of the CLIP’s visual encoder and identify the image whose embedding is the closest to the cluster center (for the label-free setting we use the pseudo-labels of the images during clustering), then we prompt LLaVA to generate attributes using this single image and the class name. Once generated, we use $k$ discriminative concepts to form the bottleneck. The second alternative is the image-wise attribute approach, where we use each image in the training set and prompt LLaVA to generate per image attributes describing the image. Once generated, we sort the concepts based on their frequency of occurrence in a class and use the most frequently occurring discriminative concepts to form the bottleneck. While the image-level concepts lead to the best coresets, it is slow and costly to prompt LLaVA to generate attributes for all the images in a large dataset such as Imagenet.
@@
Table 6: Effect of the number of per-class concepts $k$ and the method of selecting $k$ attribute-level concepts from those generated by LLaVA on the accuracy of models trained on the coresetfootnotemark: .
@@
Random concepts $48.78_{\scriptscriptstyle\pm 0.96}$ $50.15_{\scriptscriptstyle\pm 1.64}$ $50.39_{\scriptscriptstyle\pm 0.72}$
Discriminative $51.42_{\scriptscriptstyle\pm 0.18}$ $\mathbf{51.85}_{\scriptscriptstyle\pm 0.29}$ $51.22_{\scriptscriptstyle\pm 0.72}$
For CIFAR-100, this process took about nine hours to complete (in comparison CW-A can be extracted in 5 minutes without parallel computation) which is very costly compared to the small performance gains it provides over other approaches. Lastly, while the class-wise one-shot image attribute approach is better than CW-A, the additional step of clustering can be costly for larger datasets such as Imagenet. Based on these results, we used CW-A for concept generation using LLaVA.
Effect of $k$ and the method for selecting $k$ concepts. In Table 6, we show how the number of concepts extracted per class label, for creating the bottleneck in block 1 of Fig. 2 affects the selection of coresets. Once the list of class-wise attribute-level concepts is generated by LLaVA, we can select $k$ concepts per class either randomly or choose concepts unique to a class (discriminative). Our results show that using even $k=1$ is sufficient to surpass the performance using a random subsetfootnotemark: . This performance increases when we keep discriminative concepts in our concept bottleneck, with $k=5$ achieving the best results. While the size of the concept bottleneck need not be large to find good coresets, it is helpful to take a sample’s visual similarity with a set of concepts rather than a single concept per class. Thus, we used $5$ discriminative concepts per class to form the bottleneck.
Table 7: Effect of using different LLMs for generating concepts used in the concept bottleneckfootnotemark: .
LLM Accuracy
GPT-3 $51.05_{\scriptscriptstyle\pm 0.71}$
Phi-3 $51.30_{\scriptscriptstyle\pm 0.26}$
LLaVA $51.85_{\scriptscriptstyle\pm 0.29}$
Table 8: Effectfootnotemark: of using various CLIP backbones for computing the visual-concept similarity in Eq. 2.
CLIP backbone Accuracy
ViT B-16 $48.89_{\scriptscriptstyle\pm 1.01}$
ViT L-14 $49.54_{\scriptscriptstyle\pm 1.82}$
ResNet-50 $50.56_{\scriptscriptstyle\pm 1.67}$
ViT B-32 $51.85_{\scriptscriptstyle\pm 0.29}$
Table 9: Effectfootnotemark: of using different number training epochs $T$ for computing the score in Eq. 4.
$T$ Accuracy
10 $46.63_{\scriptscriptstyle\pm 0.08}$
20 $48.40_{\scriptscriptstyle\pm 0.10}$
50 $50.50_{\scriptscriptstyle\pm 0.31}$
100 $51.85_{\scriptscriptstyle\pm 0.29}$
Effect of using different LLMs/VLMs for concept extraction. We evaluated three different VLMs for concept extraction in block 1 of Fig. 2. We used GPT-3 and two open source VLMs namely Phi-3-Mini-4K-Instruct with 3.8 billion parameters, and LLaVA with 7 billion parameters. In Table 9, we find that the performance of our method remains stable regardless of the LLM used, indicating that even smaller LLMs are effective at generating concepts that produce high performing coresets.
Effect of using different VLMs for measuring visual and concept similarity. Here we evaluated different CLIP backbones to compute the similarity between the visual and concept features used in Eq. 2 (block 2 of Fig. 2). Our results in Table 9 show that our concept-based method achieves significantly better performance than Randomfootnotemark: for all the backbones with ViT B-32 backbone performing the best. Thus, we used this backbone for all the experiments in the paper.
Effect of number of epochs $T$. Here we evaluated the effect of using different number of epochs $T$ used for training the concept bottleneck layer for computing the concept-based score. Our results in Table 9 show that $T\geq 50$ is enough to achieve concept-based scores that lead to selection of high performing coresets. We used $T=100$ in our experiments since that achieves the best performance.
Evaluation on medical and affective computing tasks. In App. C.2, we present an evaluation of using our approach for CS for biomedical entity recognition (using BloodMNIST acevedo2020dataset ) and emotion recognition (using a subset of AffectNet mollahosseini2017affectnet ) task. Superior performance of our approach in Tables 10 and 11 on these tasks highlights the effectiveness of our approach for CS on diverse tasks.
@@ 6 Conclusion
CS finds representative samples from a large dataset, training models on which leads to models with accuracy similar to the models trained on the entire dataset. In this work, we proposed a scoring mechanism based on concept bottlenecks that allows us to compute the difficulty of a sample in terms of interpretable concepts. This method is independent of the downstream model and avoids training the downstream model on the full dataset even once. Our experiments show that training downstream models on coresets selected using our approach leads to better performance than random subsets and achieves accuracy similar to or better than the SOTA approaches based on training dynamics of the downstream model, for both the standard and label-free CS problem. Moreover, our score provides an intuitive explanation of a sample’s difficulty at a dataset level, independent of the downstream model.
Limitations: While we evaluated various LLMs for concept generation and showed their effectiveness for computing concept-based score, it is possible that for some classes LLMs produce very noisy and non-discriminative concepts, leading to poor concept-based scores. Thus, developing methods to prompt the LLMs to generate discriminative concepts to be used as the concept bottleneck is a crucial research direction and is left for future works. Another limitation is the dependence of our score on the VLMs for computing similarity between visual and concept features. Although we evaluated our approach on a diverse set of tasks including a biomedical task and various CLIP-based VLMs, it is possible that these similarity scores may not work well for certain tasks, warranting additional fine-tuning to make the similarity scores meaningful. Hence, in practice, additional time/compute may be required to identify or fine-tune a VLM for CS.
@@
Appendix A Additional related work
Concept-based interpretability: Interpretability methods can be broadly classified as post-hoc methods (do not impose any model constraints) or by-design methods. Post-hoc methods include Gradient-weighted Class Activation Mapping approaches [3, 59, 46, 23] that trace network gradients to identify the input areas that guide predictions and Explanation Generation methods [63, 48, 32, 20] that require models to produce explanations for visual tasks by conditioning their predictions on captioning models or incorporating visual evidence to ground explanations [21, 49]. Interpretable-by-design methods, such as Prototype methods, optimize a metric space where classifications are based on distances to class prototypes, identifying important input regions but often obscuring their semantic content [47, 9, 64, 57, 72].
Concept Bottleneck Models (CBMs) are a part of interpretable-by-design approaches that use human-understandable attributes as an intermediate layer for predictions. A recent advancement, Computational Derivation Learning (CompDL), utilizes a CBM architecture by applying a linear layer over CLIP scores between expert-designed concepts and images, improving evaluation of how well CLIP grounds concepts [84]. Post-hoc Concept Bottleneck Models (PCBMs) were recently proposed to ease the requirement of CBMs to rely on costly concept annotations and improve their accuracy compared to end-to-end models. However, PCBMs are limited by the coverage of knowledge bases, making them unsuitable for large-scale or domain-specific tasks or fine-grained classification, and their residual predictors can undermine interpretability by blending CBMs with end-to-end models.
High-level semantic-driven descriptions are also used to guide data augmentation to build an informative set [74] to make model training efficient with a good enough training set. Prior works use external knowledge bases to obtain these textual semantic concepts to guide vision models [5, 29, 54, 62]. Thus, the use of concepts has been shown to improve interpretability in various domains. However, to the best of our knowledge, we are the first ones to propose a concept-based score for the CS problem and show its competitiveness to SOTA model training dynamics-dependent approaches.
Adaptive subset selection: These works focus of improving the training convergence and efficiency of model training by selecting a new subset of data from the whole dataset, every epoch, while training a downstream model. Thus, unlike our work, these works do not prune the dataset but rather keep selecting small, potentially non-overlapping, subsets every few epochs. Both [31, 71], propose ways to select subsets every epoch or every few epochs in a way that does not use a downstream model unlike some other works such as GradMatch [30], which select subsets based on the downstream model. While our work does not target adaptive subset selection, we evaluated how well our approach performs without any change on this problem in App. C.3. Our results show that, even on for this application, our concept-based score is an effective method.
@@ Appendix B Details for concept set generation
Alternative VLMs for Concept Generation: We leverage LLaVA as our choice of VLM for concept generation, however in Table 9, we also compared against concepts extracted from GPT [79] and another smaller open source VLM (Phi-3-Mini-4K-Instruct). We see comparable performance of our CS method with concepts extracted from these models. We also experimented with retrieving concepts via another recent method, SpLiCE [4] which uses a linear optimization for sparse concept decomposition. However, a major limitation of SpLiCE is that similar to image-wise attributes as used in Table 5 it is a costly approach (SpLiCE can take up to 3 hours for $50,000$ images, significantly slower than generating class-level concepts from LLaVA).
@@ C.1 Visualizing easy/challenging samples based on concept-based score
Table 10: Performance of concept-based coreset selection for emotion recognition task on Affectnet dataset. Coresets selected by our approach results in superior performance (higher F1 score) compared to Random for standard CS.
Model Arch. Method Pruning Rates
$30\%$ $50\%$ $70\%$ $90\%$
EfficientNet Random $0.607_{\scriptscriptstyle\pm 0.006}$ $0.573_{\scriptscriptstyle\pm 0.025}$ $0.517_{\scriptscriptstyle\pm 0.006}$ $0.347_{\scriptscriptstyle\pm 0.055}$
Ours $0.603_{\scriptscriptstyle\pm 0.006}$ $0.577_{\scriptscriptstyle\pm 0.021}$ $0.537_{\scriptscriptstyle\pm 0.015}$ $0.450_{\scriptscriptstyle\pm 0.035}$
Table 11: Performance of concept-based coreset selection for biomedical image recognition task on BloodMNIST dataset. Coresets selected by our approach results in superior accuracy compared to Random for standard CS, highlighting our method’s effectiveness across various tasks.
Model Arch. Method Pruning Rates
$30\%$ $50\%$ $70\%$ $90\%$
ResNet-18 Random $94.99_{\scriptscriptstyle\pm 0.27}$ $94.79_{\scriptscriptstyle\pm 0.18}$ $91.71_{\scriptscriptstyle\pm 0.29}$ $86.50_{\scriptscriptstyle\pm 0.66}$
Ours $94.64_{\scriptscriptstyle\pm 0.60}$ $94.06_{\scriptscriptstyle\pm 0.10}$ $92.26_{\scriptscriptstyle\pm 0.19}$ $87.44_{\scriptscriptstyle\pm 1.46}$
C.2 Concept-based CS for emotion recognition and biomedical image recognition
To validate the effectiveness of our concept-based coreset selection method beyond object recognition tasks, we apply our concept-based CS approach to the task of emotion recognition and biomedical image recognition.
For emotion recognition, we use the Affectnet dataset [45] for our experiments. AffectNet is a large-scale facial expression dataset designed for training and evaluating affective computing models [73]. It contains facial images collected from the internet using web search queries for emotion-related keywords in multiple languages. Each image is manually annotated for eight discrete emotion categories: neutral, happiness, sadness, surprise, fear, disgust, anger, contempt. For our experiments, we utilize an openly available version of this dataset  [^3] https://www.kaggle.com/datasets/noamsegal/affectnet-training-data, containing roughly $16000$ training and $14000$ testing samples.
According to our approach we first use LLaVA to extract concepts for the 8 emotion classes, using the following prompt, “What are the facial features that distinguish emotion class name from other emotion types. Focus on changes in eyes, nose, lips, eyebrows, mouth. Give the output separated by commas in a line.”. We get $5-10$ distinctive facial feature concepts for every emotion, for instance for emotion class happy, we get the following concepts, “wide open eyes”, “sparking eyes”, “smiling lips”, “open mouth”, “raised eyebrows”, “flushed cheeks”, “teeth barred”. We finally select $k=5$ discriminative concepts from this list.
To test coreset performance, we use the EfficientNet model [67] and report F1 scores for our coresets in Table 10. When compared against randomly selected coresets for the various pruning ratios, coresets selected via our concept-based approach achieve better performance.
For biomedical image recognition, we use the BloodMNIST [1] dataset from the MedMNIST [80, 81] which comprises of images of normal blood cells, captured from individuals without infection, hematologic or oncologic disease and free of any pharmacologic treatment at the moment of blood collection. It consists of a total of $17,092$ images and is organized into 8 classes (basophil,eosinophil,erythroblast,immature granulocytes(myelocytes, metamyelocytes and promyelocytes),lymphocyte,monocyte,neutrophil,platelet).
Table 12: Adaptive coreset selection (Results for RBFNN and GradMatchPB are taken from tables 1 and 3 of Tukan et. al.)
Method Datasets and Pruning Rates
CIFAR-10 CIFAR-100
$70\%$ $90\%$ $70\%$ $90\%$
GradMatchPB [30] $91.89$ $90.01$ $72.57$ $60.39$
RBFNN [71] 94.4494.44{94.44}94.44 91.4091.40{91.40}91.40 $73.48$ $64.59$
Adaptive Random $93.64_{\scriptscriptstyle\pm 0.21}$ $90.49_{\scriptscriptstyle\pm 0.41}$ ${71.70}_{\scriptscriptstyle\pm 0.30}$ ${61.24}_{\scriptscriptstyle\pm 0.37}$
Ours (adaptive) $92.61_{\scriptscriptstyle\pm 0.08}$ $89.95_{\scriptscriptstyle\pm 0.26}$ ${71.74}_{\scriptscriptstyle\pm 0.33}$ $63.71_{\scriptscriptstyle\pm 0.48}$
C.3 Evaluation on adaptive subset selection
In this section, we present a comparison of using our concept-based method for the problem of adaptive subset selection [30, 31, 71] (See App. A for a description of the problem). To test our approach, we followed the experimental setup of [71], of training a ResNet-18 model on CIFAR-10/100 for $300$ epochs and changing the subset every $20$ epochs. Unlike works in the line of adaptive subset selection which focus on selecting subsets that do not overlap or train models on easy subsets first before moving to difficult ones, we simply used our CCS-based selection approach [86] to identify new subsets every $20$ epochs for this task. Due to the randomness in CCS (Alg. 1), samples being selected to form the coreset changes.
In our table 12, we compare our method to previous methods and with an adaptive random selection method (where a random subset of data is selected every $20$ epochs), which has been suggested by [31] as a strong baseline for this line of work. We observe that our method without any modifications achieves comparable performance to existing adaptive subset selection methods showing its effectiveness for this problem as well. Moreover, sampling new subsets via CCS (Alg. 1) is as efficient as sampling a random subset; thus our approach is also an efficient solution for this problem. While we believe that incorporating better subset sampling techniques as suggested by [31] may boost the performance of our method on this problem, CCS-based sampling is already quite effective.
C.4 Transferability of coresets
In this section, we present the performance of three additional downstream model architectures after training on the coresets found by our approach. Similar to the results in Table 3, our results in Table 13 show that coresets found by our approach in both standard and label-free setting achieve performance better than the random subsets on these three architectures as well. We note that the performance of the ViT model is worse than the performance of other model architectures with and without pruning (ViT-B-16 achieves an accuracy of $\approx 62\%$ compared to $\approx 78\%$ with ResNet-50 on full Imagenet dataset) due to the ViT models being data hungry in nature [14] and the fact that we used standard SGD-based training (similar to that used for training other models in the paper). We believe that using other training methods for ViTs as suggested in [69] could produce better performing ViT models.
Nonetheless, better performance than random subsets across a variety of downstream model architectures highlights the effectiveness of our approach at finding coresets without the knowledge of the architecture or training dynamics of the downstream models.
Table 13: Superior performance of downstream models with different architectures trained on our coresets for Imagenet compared to Random for both standard (Ours) and label-free (Ours-LF) CS highlight our transferability of our coresets.
Model Architecture Method Pruning Rates
$30\%$ $50\%$ $70\%$ $90\%$
MobileNet Random $62.68_{\scriptscriptstyle\pm 0.09}$ $62.49_{\scriptscriptstyle\pm 0.15}$ $61.53_{\scriptscriptstyle\pm 0.29}$ $53.80_{\scriptscriptstyle\pm 0.17}$
Ours-LF $61.60_{\scriptscriptstyle\pm 0.18}$ $61.97_{\scriptscriptstyle\pm 0.08}$ $61.73_{\scriptscriptstyle\pm 0.25}$ $54.39_{\scriptscriptstyle\pm 0.66}$
Ours $61.36_{\scriptscriptstyle\pm 0.15}$ $62.32_{\scriptscriptstyle\pm 0.22}$ $62.46_{\scriptscriptstyle\pm 0.26}$ $55.57_{\scriptscriptstyle\pm 0.13}$
RN-34 Random $73.37_{\scriptscriptstyle\pm 0.08}$ $71.71_{\scriptscriptstyle\pm 0.10}$ $67.85_{\scriptscriptstyle\pm 0.04}$ $51.29_{\scriptscriptstyle\pm 0.20}$
Ours-LF $73.61_{\scriptscriptstyle\pm 0.08}$ $71.99_{\scriptscriptstyle\pm 0.05}$ $68.42_{\scriptscriptstyle\pm 0.21}$ $53.52_{\scriptscriptstyle\pm 0.06}$
Ours $73.39_{\scriptscriptstyle\pm 0.12}$ $72.34_{\scriptscriptstyle\pm 0.13}$ $69.44_{\scriptscriptstyle\pm 0.17}$ $55.92_{\scriptscriptstyle\pm 0.02}$
ViT-B-16 Random $59.09_{\scriptscriptstyle\pm 0.49}$ $51.65_{\scriptscriptstyle\pm 0.46}$ $40.19_{\scriptscriptstyle\pm 0.13}$ $22.13_{\scriptscriptstyle\pm 0.16}$
Ours-LF $57.50_{\scriptscriptstyle\pm 0.69}$ $49.81_{\scriptscriptstyle\pm 0.89}$ $40.33_{\scriptscriptstyle\pm 0.86}$ $23.08_{\scriptscriptstyle\pm 0.27}$
Ours $57.62_{\scriptscriptstyle\pm 0.56}$ $52.67_{\scriptscriptstyle\pm 0.49}$ $42.15_{\scriptscriptstyle\pm 0.81}$ $24.43_{\scriptscriptstyle\pm 0.28}$
C.5 Algorithm for stratified sampling using CCS [86]